OpenStudy (anonymous):

The total perimeter of several rectangular shaped Olympic sized swimming pools is 750 meters. If the length of each pool is 50 meters and the width is 25 meters, how many pools are in the athletic complex?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

the perimeter of each pool is: p=2*(50+25)=...?

OpenStudy (anonymous):

150:)

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

so the requested number is: N=750/150=...?

OpenStudy (anonymous):

5

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay!!!

OpenStudy (michele_laino):

:)